# ACT-Tensor: Tensor Completion Framework for Financial Dataset Imputation

### v1

@@ Abstract
Missing data in financial panels presents a critical obstacle, undermining asset-pricing models and reducing the effectiveness of investment strategies. Such panels are often inherently multi-dimensional, spanning firms, time, and financial variables, which adds complexity to the imputation task. Conventional imputation methods often fail by flattening the data’s multidimensional structure, struggling with heterogeneous missingness patterns, or overfitting in the face of extreme data sparsity. To address these limitations, we introduce an Adaptive, Cluster-based Temporal smoothing tensor completion framework (ACT-Tensor) tailored for severely and heterogeneously missing multi-dimensional financial data panels. ACT-Tensor incorporates two key innovations: a cluster-based completion module that captures cross-sectional heterogeneity by learning group-specific latent structures; and a temporal smoothing module that proactively removes short-lived noise while preserving slow-moving fundamental trends. Extensive experiments show that ACT-Tensor consistently outperforms state-of-the-art benchmarks in terms of imputation accuracy across a range of missing data regimes, including extreme sparsity scenarios. To assess its practical financial utility, we evaluate the imputed data with an asset-pricing pipeline tailored for tensor-structured financial data. Results show that ACT-Tensor not only reduces pricing errors but also significantly improves risk-adjusted returns of the constructed portfolio. These findings confirm that our method delivers highly accurate and informative imputations, offering substantial value for financial decision-making.
@@
Keywords Asset Pricing, Missing Data Imputation, Tensor Completion, Factor Models
1 Introduction
@@
Financial data—particularly firm-level characteristic data—are widely used by researchers and practitioners in empirical asset pricing to explain and predict the cross-section of expected returns [1, 2, 3, 4] and to support the construction of systematic investment strategies [5]. These data, indexed by firm, characteristic, and time, naturally form a multi-dimensional structure, which not only captures temporal dynamics but also preserves cross-sectional heterogeneity across firms. The financial utility of asset pricing models critically depends on the completeness and quality of these multi-dimensional panels, as reliable inputs are essential for effective investment decisions.
@@
However, in practice, such fully observed firm-level characteristic panels are not always available. For example, CRSP/Compustat data show that at any given month, more than 70% of listed equities lack at least one characteristic, accounting for roughly half of the market’s total capitalization [6]. Moreover, this missingness is systematic: smaller, younger, or financially distressed firms are disproportionately incomplete [7, 8]. When these extensive, non-random missing data are handled with overly simplistic or inadequate methods, they embed systematic bias into asset pricing models, which leads to significant pricing errors and sub-optimal investment strategies.
@@
Early solutions, such as cross-sectional median filling [9, 10, 11] and discarding incomplete observations [12, 13, 14] either attenuate the signal or induce selection bias. Nevertheless, more refined techniques offer only partial relief. The generalized method of moments (GMM) framework [7] relies on a missing-at-random (MAR) assumption, while Expectation-Maximization (EM) algorithms [15, 16, 17] fill the gaps under even stricter and rarely satisfied assumptions, such as joint-normality. Matrix-based imputations [6] flatten the three-dimensional panel into a two-dimensional matrix, discarding the time dimension and the temporal dependence essential to financial characteristics.
@@
More recent work preserves the panel’s three-way structure by adapting tensor completion algorithms [18, 19, 20], while several key challenges remain unresolved. First, extreme sparsity of data breeds over-fitting, and theory shows reconstruction error grows unbounded as data coverage shrinks [21, 22, 23]. Second, the single latent factor structure applied to all firms ignores the firm-level heterogeneity [7, 6]. Most models treat the time dimension as a conventional variable [19, 20], neglecting its unique properties. Hence, they often fail to account for the non-stationarity inherent in firm characteristics.
@@
To address the challenges posed by missing data in financial panels, we propose an Adaptive, Cluster-based Temporal smoothing tensor completion framework (ACT-Tensor) designed to handle sparse, multidimensional financial datasets. Unlike conventional methods, which either discard the rich cross-sectional and temporal dependencies or fail to properly address extreme sparsity, ACT-Tensor is specifically built to preserve the multidimensional structure of the data, ensuring more stable factor estimates and accurate return forecasts. In general, our main contributions are threefold.
@@
First, we enhance tensor completion with two innovative modules: (i) a cluster-based completion module that captures the rich cross-sectional heterogeneity to address overfitting under extreme sparsity and avoid model bias in methods that apply a single global structure; and (ii) a temporal smoothing module that filters short-lived noise while preserving persistent signals to ensure imputed values reflect long-term trends and remain robust to non-stationarity. Together, they overcome the limitations of existing methods that fail to capture firm-level heterogeneity and temporal dependencies.
@@ 2 Related Work
Early Approaches. Early solutions for missing data involved filling missing values with cross-sectional medians [9, 10, 11] or confining analysis to the minority of firms with fully observed data [13, 12, 14]. While straightforward to implement, empirical evidence has shown that these ad-hoc strategies introduce significant bias and undermine the robustness of model estimations [9].
@@
Matrix Factorization Based Models. To exploit the cross-sectional correlations, many approaches collapse the three-way panel into matrices and applying principal-component analysis [24, 25] or matrix completion methods [16, 6]. While this flatten-and-factor strategy improves upon simple heuristics, it applies a single low-rank structure on all firms and discard both temporal dependencies [26] and firm-level heterogeneous effects [19].
@@
Estimation Based Models. Another line of methods integrate imputation directly into model estimation. GMM-based methods [7] require a MAR assumption, where the probability that a characteristic is missing is independent of its unobserved value. Similarly, EM-based techniques [15, 16, 17] rely on strong distributional assumptions, such as joint normality, and impute each period in isolation, discarding valuable time-series information. Consequently, both families of methods can produce biased or unstable estimates when data are sparse.
@@
Tensor Completion Based Models. A body of work preserves the panel’s three-way structure by applying low-rank tensor decompositions for imputation [19, 20, 18]. Empirically, these tensor methods outperform matrix-based approaches, especially when data are sparse [18, 27]. However, key limitations remain: (i) extreme sparsity still leads to overfitting, since learning a single global rank from sparse observations leads to reconstruction error to explode as data-density shrinks [21, 22, 23]; (ii) the assumption of one latent structure for all firms ignores cross-sectional heterogeneity in size, industry, and life cycle [7, 6]; and (iii) most tensor-completion models treat the time dimension as a conventional variable and ignore its unique properties. Even time-variant extensions [28, 29] are computationally burdensome and untested on large financial panels with extreme sparsity.
@@
Machine Learning Based Models. Recent studies also apply deep learning techniques, including recurrent neural networks [30, 31] and Transformers [32, 33], to impute missing financial data. Although these architectures can model complex nonlinearities, they treat each firm’s observations as plain vectors or sequences, overlooking the rich cross-sectional and characteristic-level interactions captured by a multidimensional tensor. Moreover, deep learning models often act as black boxes with limited interpretability and theoretical guarantees, and they demand extensive hyperparameter tuning and regularization to avoid overfitting. Consequently, purely ML-based imputations struggle to deliver consistently high accuracy without embedding domain-specific structure.
@@
In summary, we overcome the challenges faced by existing methods through introducing ACT-Tensor with two innovative modules to (i) handle extreme data sparsity without over-fitting, (ii) capture firm-level heterogeneity via cluster-based completion, and (iii) accommodate temporal non-stationarity with data-driven smoothing. Unlike prior studies that judge success only by statistical error, we also evaluate imputations by their impact on pricing accuracy and return forecasts. The comprehensive result shows that ACT-Tensor delivers state-of-the-art recovery of missing data and significantly improves pricing accuracy, thereby demonstrating the practical value of a unified approach that rigorously connects imputation advances to asset-pricing performance.
3 ACT-Tensor Imputation Framework
@@
Missing data in firm-characteristic panels undermines factor estimates and return forecasts in asset-pricing models. To counter this problem and overcome the shortcomings of existing imputation methods, we introduce ACT-Tensor, which divides the problem into two innovative, complementary modules: (i) cluster-based completion (Section 3.2) and (ii) temporal smoothing (Section 3.3).
3.1 Tensor Factorization and Completion
@@
We treat the firm-characteristic panel as a third-order tensor $\mathcal{X}\in\mathbb{R}^{T\times N\times L}$, where $T$ is the number of months, $N$ is the number of firms, and $L$ is the number of characteristics. To exploit the low-rank structure of $\mathcal{X}$, we adopt the CANDECOMP/PARAFAC (CP) decomposition. For a chosen rank $R$, the completed tensor $\hat{\mathcal{X}}$ is written in terms of three loading matrices $U\in\mathbb{R}^{T\times R}$, $V\in\mathbb{R}^{N\times R}$, and $W\in\mathbb{R}^{L\times R}$ with component weights $\boldsymbol{\gamma}=(\gamma_{1},\dots,\gamma_{R})^{\intercal}$:
$\hat{\mathcal{X}}=[\![\boldsymbol{\gamma}\,;\,U,\,V,\,W\,]\!]=\sum_{r=1}^{R}\gamma_{r}\,\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}.$ (1)
@@
To remove scale ambiguity we set $\gamma_{r}=1$ for $r=1,\dots,R$ [34]. Each entry of $\hat{\mathcal{X}}$ is simply:
$\hat{x}_{t,n,\ell}=\sum_{r=1}^{R}\gamma_{r}\,U_{tr}\,V_{nr}\,W_{\ell r},$ (2)
[…]
$\min_{U,V,W}\left\|\mathcal{P}_{\Omega}(\mathcal{X})-\mathcal{P}_{\Omega}\left([\![U,V,W]\!]\right)\right\|_{F}^{2}+\lambda\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}+\|W\|_{F}^{2}\right),$ (3)
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm, $\mathcal{P}_{\Omega}$ masks unobserved entries and $\lambda$ is an $\ell_{2}$ regularizer that guards against over-fitting.
@@
Although ACT-Tensor is fully compatible with Tucker or Tensor-Train decompositions, we focus on CP here for its computational simplicity and its proven effectiveness in large, sparse settings [19, 18], avoiding the dense cores or multiple-core overhead required by alternative tensor decompositions.
3.2 Cluster-based Completion
@@ 3.2.1 Clustering of Companies
$\mathbf{X}_{n}=\mathcal{X}\left({:,n,:}\right)\in\mathbb{R}^{T\times L},\quad\mathbf{v}_{n}=\operatorname{vec}\!\left(\mathbf{X}_{n}\right)\in\mathbb{R}^{TL},$ (4)
@@
$\min_{z\in\{1,\dots,K\}^{N},\,\{\boldsymbol{\mu}_{k}\}_{k=1}^{K}}\,\sum_{n=1}^{N}\left\lVert\mathbf{v}_{n}-\boldsymbol{\mu}_{\,z_{n}}\right\rVert_{2}^{2},$ (5)
where $z_{n}$ is firm $n$’s cluster label and $\mu_{k}\in\mathbb{R}^{TL}$ is the $k$th centroid. Let $\mathcal{I}_{k}=\{n:z_{n}=k\}$ be the firms in cluster $k$, and define its observed-entry ratio as:
$\rho_{k}=\frac{1}{|\mathcal{I}_{k}|\,T\,L}\sum_{n\in\mathcal{I}_{k}}\sum_{t=1}^{T}\sum_{\ell=1}^{L}\mathbf{1}\left[(t,n,\ell)\in\Omega\right],$ (6)
@@
$\mathcal{X}_{k}=\mathcal{X}_{:\,,\,\mathcal{I}_{k}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{k}|\times L}.$ (7)
Given a threshold $\tau$, we call cluster $k$ dense if $\rho_{k}\geq\tau$ and sparse otherwise. $\tau=40\%$ is chosen to ensure a stable completion [22, 18, 35]. Thus, after density–driven clustering, we categorize the clusters into dense $\mathcal{D}=\{d:\rho_{d}\geq\tau\}$ and sparse $\mathcal{S}=\{s:\rho_{s}<\tau\}$.
3.2.2 Cluster-based Completion

### v3

@@ Abstract.
Missing data in financial panels presents a critical obstacle, undermining asset-pricing models and reducing the effectiveness of investment strategies. Such panels are often inherently multi-dimensional, spanning firms, time, and financial variables, which adds complexity to the imputation task. Conventional imputation methods often fail by flattening the data’s multidimensional structure, struggling with heterogeneous missingness patterns, or overfitting in the face of extreme data sparsity. To address these limitations, we introduce an Adaptive, Cluster-based Temporal smoothing tensor completion framework (ACT‑Tensor) tailored for severely and heterogeneously missing multi-dimensional financial data panels. ACT-Tensor incorporates two key innovations: a cluster‑based completion module that captures cross‑sectional heterogeneity by learning group-specific latent structures; and a temporal smoothing module that proactively removes short‑lived noise while preserving slow‑moving fundamental trends. Extensive experiments show that ACT‑Tensor consistently outperforms state-of-the-art benchmarks in terms of imputation accuracy across a range of missing data regimes, including extreme sparsity scenarios. To assess its practical financial utility, we evaluate the imputed data with an asset‐pricing pipeline tailored for tensor-structured financial data. Results show that ACT-Tensor not only reduces pricing errors but also significantly improves risk-adjusted returns of the constructed portfolio. These findings confirm that our method delivers highly accurate and informative imputations, offering substantial value for financial decision-making.
@@
Asset Pricing, Missing Data Imputation, Tensor Completion, Factor Models
journalyear: 2025copyright: acmlicensedconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770408isbn: 979-8-4007-2220-2/2025/11ccs: Applied computing Operations researchccs: Computing methodologies Artificial Intelligence
1. Introduction
@@
Financial data—particularly firm-level characteristic data—are widely used by researchers and practitioners in empirical asset pricing to explain and predict the cross-section of expected returns (Fama and French, 1992, 1993; Carhart, 1997; Pástor and Stambaugh, 2003) and to support the construction of systematic investment strategies (Asness et al., 2019). These data, indexed by firm, characteristic, and time, naturally form a multi-dimensional structure, which not only captures temporal dynamics but also preserves cross-sectional heterogeneity across firms. The financial utility of asset pricing models critically depends on the completeness and quality of these multi-dimensional panels, as reliable inputs are essential for effective investment decisions.
@@
However, in practice, such fully observed firm‑level characteristic panels are not always available. For example, CRSP/Compustat data show that at any given month, more than 70% of listed equities lack at least one characteristic, accounting for roughly half of the market’s total capitalization (Bryzgalova et al., 2025). Moreover, this missingness is systematic: smaller, younger, or financially distressed firms are disproportionately incomplete (Freyberger et al., 2025; Easton et al., 2020). When these extensive, non‑random missing data are handled with overly simplistic or inadequate methods, they embed systematic bias into asset pricing models, which leads to significant pricing errors and sub-optimal investment strategies.
@@
Early solutions, such as cross-sectional median filling (Gu et al., 2020; Kozak et al., 2020; Light et al., 2017) and discarding incomplete observations (Freyberger et al., 2020; Lewellen, 2014; Kelly et al., 2019) either attenuate the signal or induce selection bias. Nevertheless, more refined techniques offer only partial relief. The generalized method of moments (GMM) framework (Freyberger et al., 2025) relies on a missing-at-random (MAR) assumption, while Expectation-Maximization (EM) algorithms (Chen and McCoy, 2022; Jin et al., 2021; Bańbura and Modugno, 2014) fill the gaps under even stricter and rarely satisfied assumptions, such as joint-normality. Matrix‑based imputations (Bryzgalova et al., 2025) flatten the three‑dimensional panel into a two‑dimensional matrix, discarding the time dimension and the temporal dependence essential to financial characteristics.
@@
More recent work preserves the panel’s three-way structure by adapting tensor completion algorithms (Zhou et al., 2023; Acar et al., 2011; Choi et al., 2019), while several key challenges remain unresolved. First, extreme sparsity of data breeds over‑fitting, and theory shows reconstruction error grows unbounded as data coverage shrinks (Montanari and Sun, 2018; Mu et al., 2014; Liu and Moitra, 2020). Second, the single latent factor structure applied to all firms ignores the firm-level heterogeneity (Freyberger et al., 2025; Bryzgalova et al., 2025). Most models treat the time dimension as a conventional variable (Acar et al., 2011; Choi et al., 2019), neglecting its unique properties. Hence, they often fail to account for the non-stationarity inherent in firm characteristics.
@@
To address the challenges posed by missing data in financial panels, we propose an Adaptive, Cluster-based Temporal smoothing tensor completion framework (ACT-Tensor) designed to handle sparse, multidimensional financial datasets. Unlike conventional methods, which either discard the rich cross‑sectional and temporal dependencies or fail to properly address extreme sparsity, ACT‑Tensor is specifically built to preserve the multidimensional structure of the data, ensuring more stable factor estimates and accurate return forecasts. In general, our main contributions are threefold.
@@
First, we enhance tensor completion with two innovative modules: (i) a cluster‑based completion module that captures the rich cross‑sectional heterogeneity to address overfitting under extreme sparsity and avoid model bias in methods that apply a single global structure; and (ii) a temporal smoothing module that filters short‑lived noise while preserving persistent signals to ensure imputed values reflect long‑term trends and remain robust to non-stationarity. Together, they overcome the limitations of existing methods that fail to capture firm-level heterogeneity and temporal dependencies.
@@ 2. Related Work
Early Approaches. Early solutions for missing data involved filling missing values with cross-sectional medians (Gu et al., 2020; Kozak et al., 2020; Light et al., 2017) or confining analysis to the minority of firms with fully observed data (Lewellen, 2014; Freyberger et al., 2020; Kelly et al., 2019). While straightforward to implement, empirical evidence has shown that these ad-hoc strategies introduce significant bias and undermine the robustness of model estimations (Gu et al., 2020).
@@
Matrix Factorization Based Models. To exploit the cross-sectional correlations, many approaches collapse the three-way panel into matrices and applying principal‑component analysis (Bai and Ng, 2017; Cahan et al., 2023) or matrix completion methods (Jin et al., 2021; Bryzgalova et al., 2025). While this flatten-and-factor strategy improves upon simple heuristics, it applies a single low‑rank structure on all firms and discard both temporal dependencies (Lettau, 2024) and firm-level heterogeneous effects (Acar et al., 2011).
@@
Estimation Based Models. Another line of methods integrate imputation directly into model estimation. GMM‑based methods (Freyberger et al., 2025) require a MAR assumption, where the probability that a characteristic is missing is independent of its unobserved value. Similarly, EM-based techniques (Chen and McCoy, 2022; Jin et al., 2021; Bańbura and Modugno, 2014) rely on strong distributional assumptions, such as joint normality, and impute each period in isolation, discarding valuable time-series information. Consequently, both families of methods can produce biased or unstable estimates when data are sparse.
@@
Tensor Completion Based Models. A body of work preserves the panel’s three-way structure by applying low-rank tensor decompositions for imputation (Acar et al., 2011; Choi et al., 2019; Zhou et al., 2023). Empirically, these tensor methods outperform matrix-based approaches, especially when data are sparse (Zhou et al., 2023; Han et al., 2024). However, key limitations remain: (i) extreme sparsity still leads to overfitting, since learning a single global rank from sparse observations leads to reconstruction error to explode as data-density shrinks (Montanari and Sun, 2018; Mu et al., 2014; Liu and Moitra, 2020); (ii) the assumption of one latent structure for all firms ignores cross-sectional heterogeneity in size, industry, and life cycle (Freyberger et al., 2025; Bryzgalova et al., 2025); and (iii) most tensor-completion models treat the time dimension as a conventional variable and ignore its unique properties. Even time-variant extensions (Sun et al., 2020; Mardani et al., 2015) are computationally burdensome and untested on large financial panels with extreme sparsity.
@@
Machine Learning Based Models. Recent studies also apply deep learning techniques, including recurrent neural networks (Che et al., 2018; Cao et al., 2018) and Transformers (Beckmeyer and Wiedemann, 2022; Du et al., 2023), to impute missing financial data. Although these architectures can model complex nonlinearities, they treat each firm’s observations as plain vectors or sequences, overlooking the rich cross-sectional and characteristic-level interactions captured by a multidimensional tensor. Moreover, deep learning models often act as black boxes with limited interpretability and theoretical guarantees, and they demand extensive hyperparameter tuning and regularization to avoid overfitting. Consequently, purely ML-based imputations struggle to deliver consistently high accuracy without embedding domain-specific structure.
Advances in Tensor and Matrix Learning. Recent progress in high-dimensional matrix and tensor learning has advanced modeling, inference, and prediction for multi-way data. In matrix-variate settings, constrained factor models capture cross-sectional and temporal dependence (Chen et al., 2019), threshold structures handle regime shifts (Liu and Chen, 2022), and dynamic matrix-factor approaches extend to transport and spatial-temporal systems (Chen and Chen, 2022; Chen et al., 2020). Inference for high-dimensional matrix factors (Chen and Fan, 2023) and factor-augmented regression (Chen et al., 2024a) further enhance interpretability and robustness.
For tensors, low-rank CP/Tucker decompositions exploit multi-modal structure for efficient estimation. Semiparametric tensor factor analysis via iteratively projected SVD attains rate-optimal recovery (Chen et al., 2024d; CHEN et al., 2020), and distributed tensor PCA addresses heterogeneous data across sites (Chen et al., 2025). Supervised tensor methods extend to discriminant analysis and classification with CP-structured low-rank forms, including incomplete tensors (Chen et al., 2024b, c). Tensor structures have also been embedded into deep models such as tensor-view topological GNNs (Wen et al., 2024), tensor-fused graph contrastive learning (Wu et al., 2025), tensor-augmented Transformers (Kong et al., 2024), and tensorized uncertainty quantification (Wu et al., 2024a, b). Building on this literature, our work adapts tensor learning to large, sparsely observed financial panels. ACT-Tensor integrates low-rank estimation, distributed learning, and temporal regularization, bridging statistical tensor modeling with practical financial imputation and pricing applications.
@@
In summary, we overcome the challenges faced by existing methods through introducing ACT‑Tensor with two innovative modules to (i) handle extreme data sparsity without over‑fitting, (ii) capture firm‑level heterogeneity via cluster‑based completion, and (iii) accommodate temporal non‑stationarity with data‑driven smoothing. Unlike prior studies that judge success only by statistical error, we also evaluate imputations by their impact on pricing accuracy and return forecasts. The comprehensive result shows that ACT‑Tensor delivers state‑of‑the‑art recovery of missing data and significantly improves pricing accuracy, thereby demonstrating the practical value of a unified approach that rigorously connects imputation advances to asset‑pricing performance.
3. ACT-Tensor Imputation Framework
@@
Missing data in firm‑characteristic panels undermines factor estimates and return forecasts in asset‑pricing models. To counter this problem and overcome the shortcomings of existing imputation methods, we introduce ACT-Tensor, which divides the problem into two innovative, complementary modules: (i) cluster-based completion (Section 3.2) and (ii) temporal smoothing (Section 3.3).
3.1. Tensor Factorization and Completion
@@
We treat the firm-characteristic panel as a third‐order tensor $\mathcal{X}\in\mathbb{R}^{T\times N\times L}$, where $T$ is the number of months, $N$ is the number of firms, and $L$ is the number of characteristics. To exploit the low-rank structure of $\mathcal{X}$, we adopt the CANDECOMP/PARAFAC (CP) decomposition. For a chosen rank $R$, the completed tensor $\hat{\mathcal{X}}$ is written in terms of three loading matrices $U\in\mathbb{R}^{T\times R}$, $V\in\mathbb{R}^{N\times R}$, and $W\in\mathbb{R}^{L\times R}$ with component weights $\boldsymbol{\gamma}=(\gamma_{1},\dots,\gamma_{R})^{\intercal}$:
(1) $\hat{\mathcal{X}}=[\![\boldsymbol{\gamma}\,;\,U,\,V,\,W\,]\!]=\sum_{r=1}^{R}\gamma_{r}\,\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}.$
@@
To remove scale ambiguity we set $\gamma_{r}=1$ for $r=1,\dots,R$ (Kolda and Bader, 2009). Each entry of $\hat{\mathcal{X}}$ is simply:
(2) $\hat{x}_{t,n,\ell}=\sum_{r=1}^{R}\gamma_{r}\,U_{tr}\,V_{nr}\,W_{\ell r},$
where $\mathbf{u}_{r}\in\mathbb{R}^{T}$, $\mathbf{v}_{r}\in\mathbb{R}^{N}$, and $\mathbf{w}_{r}\in\mathbb{R}^{L}$ are the $r$th columns of $U$, $V$, and $W$. When the original tensor $\mathcal{X}$ contains missing values, we estimate $(U,V,W)$ by minimizing the reconstruction error over the observed set $\Omega$:
(3) $\min_{U,V,W}\left\|\mathcal{P}_{\Omega}(\mathcal{X})-\mathcal{P}_{\Omega}\left([\![U,V,W]\!]\right)\right\|_{F}^{2}+\lambda\left(\|U\|_{F}^{2}+\|V\|_{F}^{2}+\|W\|_{F}^{2}\right),$
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm, $\mathcal{P}_{\Omega}$ masks unobserved entries and $\lambda$ is an $\ell_{2}$ regularizer that guards against over‑fitting.
@@
Although ACT-Tensor is fully compatible with Tucker or Tensor‐Train decompositions, we focus on CP here for its computational simplicity and its proven effectiveness in large, sparse settings (Acar et al., 2011; Zhou et al., 2023), avoiding the dense cores or multiple‐core overhead required by alternative tensor decompositions.
3.2. Cluster-based Completion
@@ 3.2.1. Clustering of Companies
(4) $\mathbf{X}_{n}=\mathcal{X}\left({:,n,:}\right)\in\mathbb{R}^{T\times L},\quad\mathbf{v}_{n}=\operatorname{vec}\!\left(\mathbf{X}_{n}\right)\in\mathbb{R}^{TL},$
@@
(5) $\min_{z\in\{1,\dots,K\}^{N},\,\{\boldsymbol{\mu}_{k}\}_{k=1}^{K}}\,\sum_{n=1}^{N}\left\lVert\mathbf{v}_{n}-\boldsymbol{\mu}_{\,z_{n}}\right\rVert_{2}^{2},$
where $z_{n}$ is firm $n$’s cluster label and $\mu_{k}\in\mathbb{R}^{TL}$ is the $k$th centroid. Let $\mathcal{I}_{k}=\{n:z_{n}=k\}$ be the firms in cluster $k$, and define its observed‐entry ratio as:
(6) $\rho_{k}=\frac{1}{|\mathcal{I}_{k}|\,T\,L}\sum_{n\in\mathcal{I}_{k}}\sum_{t=1}^{T}\sum_{\ell=1}^{L}\mathbf{1}\left[(t,n,\ell)\in\Omega\right],$
@@
(7) $\mathcal{X}_{k}=\mathcal{X}_{:\,,\,\mathcal{I}_{k}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{k}|\times L}.$
Given a threshold $\tau$, we call cluster $k$ dense if $\rho_{k}\geq\tau$ and sparse otherwise. $\tau=40\%$ is chosen to ensure a stable completion (Mu et al., 2014; Zhou et al., 2023; Jain and Oh, 2014). Thus, after density–driven clustering, we categorize the clusters into dense $\mathcal{D}=\{d:\rho_{d}\geq\tau\}$ and sparse $\mathcal{S}=\{s:\rho_{s}<\tau\}$.
3.2.2. Cluster-based Completion